November 1996  ICTCM IX (Reno)  3

Example: The Parachute Problem

Nagle and Saff (Addison–Wesley, 1996, p. 112)

A parachutist whose mass is 75 kg drops from a helicopter hovering 4000 m above the ground and falls toward the earth under the influence of gravity. Assume the gravitational force is constant and that the force due to air resistance is proportional to the velocity of the parachutist, with a proportionality constant k_{1} = 15 kg/sec when the chute is closed and with a constant k_{2} = 105 kg/sec when the chute is open. If the chute does not open until 1 min after the parachutist leaves the helicopter, after how many seconds will she hit the ground?